# A mixed finite element method for a class of fourth-order stochastic evolution equations with multiplicative noise

### v1

@@ 3. A fully-discrete mixed finite element method
$\quad+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\xi}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right].$
For the terms $I_{6}$ to $I_{8}$, we note Lemma 3.2 and apply similar argument to obtain
@@
$\quad+Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\|{\nabla\boldsymbol{u}_{h}^{n}}\right\|_{\mathbb{L}^{2}}^{2}\left\|{\boldsymbol{u}(s)-\boldsymbol{u}(t_{n})}\right\|_{\mathbb{L}^{\infty}}^{2}\mathrm{d}s\right]$
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{H}^{1}}^{2}\right]+Ch^{4}+C\kappa k^{1-\delta},$
$\left|{I_{8}}\right|$ $\leq C\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\int_{t_{\ell-1}}^{t_{\ell}}\left\|{\boldsymbol{u}(s)}\right\|_{\mathbb{L}^{\infty}}^{2}\mathrm{d}s\right]+\epsilon h^{2}\mathbb{E}\left[\left\|{\boldsymbol{u}}\right\|_{L^{\infty}(\mathbb{L}^{\infty})}^{2}\right]$
$\quad+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon\mathbb{E}\left[\mathds{1}_{\Omega_{\kappa,n}}\int_{t_{n-1}}^{t_{n}}\left\|{\nabla\boldsymbol{u}(s)-\nabla\boldsymbol{u}(t_{n})}\right\|_{\mathbb{L}^{2}}^{2}\mathrm{d}s\right]$
$\leq C\kappa k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ch^{2}+\epsilon k^{1-\delta}+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right].$
@@
$\leq\mathbb{E}\left[\left\|{\boldsymbol{\theta}^{0}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ch^{2}+C(1+\kappa)k^{\frac{1}{2}-\delta}+C(1+\kappa^{2})k\sum_{\ell=1}^{n}\mathbb{E}\left[\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right].$ (3.41)
@@
$=:I_{1}+I_{2}+\cdots+I_{13}.$ (3.44)
We will estimate each term on the last line. In the following, whenever appropriate, we always use Proposition 2.2 and 3.5 without further mention. Let $\epsilon>0$. For the first term, by (2.16) and Young’s inequality we have
@@
$\leq Ce^{C\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right)+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right].$
@@
$\leq Cnh^{4}k^{-1}+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right].$
Similarly, noting Proposition 2.2 we have
@@
$\leq Ce^{C\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right)+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right],$
@@
Next, for the term $I_{6}$ we apply Young’s inequality, (2.10), (2), and (2.13). By Proposition 2.2 (Hölder continuity of $\boldsymbol{u}$) and the Sobolev embeddings, we obtain for any $\delta>0$,
@@
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$
@@
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+C(1+\kappa^{4})e^{C\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right),$ (3.46)
@@
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{4}}^{2}\right]+Ch^{4}\mathbb{E}\left[\left\|{\boldsymbol{H}}\right\|_{L^{\infty}_{T}(\mathbb{L}^{2})}^{2}\right]$
$\quad+Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\sup_{t\in[t_{\ell-1},t_{\ell}]}\left\|{\boldsymbol{H}(t)}\right\|_{\mathbb{L}^{2}}^{2}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{4}}^{2}\right]$
@@
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+C(1+\kappa)e^{C\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right),$
@@
where in the last step we also used (2.16) and (2.22), Proposition 3.5, and the Hölder continuity of $\boldsymbol{u}$. Next, similarly we have
@@
$\left|{I_{8}}\right|$ $\leq\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\|{\boldsymbol{u}_{h}^{\ell}}\right\|_{\mathbb{L}^{4}}\big{(}\left\|{\boldsymbol{\eta}^{\ell}}\right\|_{\mathbb{L}^{2}}+\left\|{\boldsymbol{\xi}^{\ell}}\right\|_{\mathbb{L}^{2}}+\left\|{\boldsymbol{H}(s)-\boldsymbol{H}(t_{\ell})}\right\|_{\mathbb{L}^{2}}\big{)}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{4}}\right]\mathrm{d}s$
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{4}}^{2}\right]+C\mathbb{E}\left[\max_{j\leq n}\left\|{\boldsymbol{\eta}^{j}}\right\|_{\mathbb{L}^{2}}^{4}\right]^{\frac{1}{2}}\mathbb{E}\left[\left(k\sum_{\ell=1}^{n}\left\|{\boldsymbol{u}_{h}^{\ell}}\right\|_{\mathbb{H}^{1}}^{2}\right)^{2}\right]^{\frac{1}{2}}$
@@
$\quad+C\kappa k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\xi}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+C\kappa\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\|{\boldsymbol{H}(s)-\boldsymbol{H}(t_{\ell})}\right\|_{\mathbb{L}^{2}}^{2}\mathrm{d}s\right]$

### v2

@@ 3. A fully-discrete mixed finite element method
$\leq Ch^{2}\mathbb{E}\left[\left\|{\boldsymbol{u}}\right\|_{L^{\infty}_{T}(\mathbb{H}^{2})}^{4}+\left(k\sum_{\ell=1}^{n}\left\|{\boldsymbol{H}_{h}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right)^{2}\right]+Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$
$\quad+Ck^{1-\delta}\mathbb{E}\left[\left\|{\boldsymbol{u}}\right\|_{\mathcal{C}^{1/2-\delta}_{T}(\mathbb{H}^{2})}^{4}+\left(k\sum_{\ell=1}^{n}\left\|{\boldsymbol{H}_{h}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right)^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$
$\leq Ch^{2}+{Ck^{1-\delta}}+Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right],$
where in the last step we also utilised (3.2). Next, with the aim of estimating $I_{5}$, we write $\boldsymbol{u}(s)=\big(\boldsymbol{u}(s)-\boldsymbol{u}(t_{\ell-1})\big)+\boldsymbol{u}(t_{\ell-1})$. By the triangle inequality, employing Hölder continuity in time for $\boldsymbol{u}$, we infer that for $s\in(t_{\ell-1},t_{\ell})$,
$\left\|{\boldsymbol{u}(s)}\right\|_{\mathbb{L}^{\infty}}\leq k^{\frac{1}{2}(1-\delta)}\left\|{\boldsymbol{u}}\right\|_{\mathcal{C}^{1/2-\delta}_{T}(\mathbb{H}^{2})}+\left\|{\boldsymbol{u}(t_{\ell-1})}\right\|_{\mathbb{L}^{\infty}}.$
Therefore, by Hölder’s and Young’s inequalities and (2.22), for any $\epsilon>0$ we have
$\left|{I_{5}}\right|$ $\leq\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\|{\boldsymbol{u}(s)}\right\|_{\mathbb{L}^{\infty}}\left(\left\|{\boldsymbol{\eta}^{\ell}}\right\|_{\mathbb{L}^{2}}+\left\|{\boldsymbol{\xi}^{\ell}}\right\|_{\mathbb{L}^{2}}+\left\|{\boldsymbol{H}(s)-\boldsymbol{H}(t_{\ell})}\right\|_{\mathbb{L}^{2}}\right)\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}\mathrm{d}s\right]$
$\leq{Ck^{1-\delta}\mathbb{E}\left[\left\|{\boldsymbol{u}}\right\|_{\mathcal{C}^{1/2-\delta}_{T}(\mathbb{H}^{2})}^{2}\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}k\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{u}(t_{\ell-1})}\right\|_{\mathbb{L}^{\infty}}^{2}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]}$
$\quad+Ch^{4}+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\xi}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\|{\boldsymbol{H}(s)-\boldsymbol{H}(t_{\ell})}\right\|_{\mathbb{L}^{2}}^{2}\mathrm{d}s\right]$
$\leq Ch^{4}+Ck^{1-\delta}\mathbb{E}\left[\left\|{\boldsymbol{u}}\right\|_{\mathcal{C}^{1/2-\delta}_{T}(\mathbb{H}^{2})}^{8}+\left(k\sum_{\ell=1}^{n}\left\|{\boldsymbol{u}_{h}^{\ell}}\right\|_{\mathbb{H}^{1}}^{2}\right)^{2}\right]+C\kappa k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$
$\quad+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\xi}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$
$\leq Ch^{4}+Ck^{1-\delta}+C\kappa k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\xi}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right],$
where in the penultimate step we used the boundedness of the Ritz projection in $\mathbb{H}^{1}$ and the regularity of $\boldsymbol{u}$, while in the last step we utilised (3.2). For the terms $I_{6}$ and $I_{7}$, we note Lemma 3.2 and apply similar argument to obtain
@@
$\quad+Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\|{\nabla\boldsymbol{u}_{h}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\left\|{\boldsymbol{u}(s)-\boldsymbol{u}(t_{\ell})}\right\|_{\mathbb{L}^{\infty}}^{2}\mathrm{d}s\right]$
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{H}^{1}}^{2}\right]+Ch^{2}+Ck^{1-\delta}\mathbb{E}\left[\left\|{\boldsymbol{u}}\right\|_{\mathcal{C}^{\frac{1}{2}-\delta}_{T}(\mathbb{H}^{2})}^{2}\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}k\left\|{\nabla\boldsymbol{u}_{h}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{H}^{1}}^{2}\right]+Ch^{2}+Ck^{1-\delta}\mathbb{E}\left[\left\|{\boldsymbol{u}}\right\|_{\mathcal{C}^{\frac{1}{2}-\delta}_{T}(\mathbb{H}^{2})}^{4}+\left(\sum_{\ell=1}^{n}k\left\|{\nabla\boldsymbol{u}_{h}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right)^{2}\right]$
$\leq Ch^{2}+Ck^{1-\delta}+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{H}^{1}}^{2}\right].$
For $I_{8}$, we proceed as in the estimate for $I_{5}$ to obtain
$\left|{I_{8}}\right|$ $\leq{C\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\|{\boldsymbol{u}(s)}\right\|_{\mathbb{L}^{\infty}}\left(\left\|{\nabla\boldsymbol{\rho}^{\ell}}\right\|_{\mathbb{L}^{2}}+\left\|{\nabla\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}+\left\|{\boldsymbol{u}(s)-\boldsymbol{u}(t_{\ell})}\right\|_{\mathbb{H}^{1}}\right)\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}\mathrm{d}s\right]}$
$\leq Ch^{2}+Ck^{1-\delta}+C\kappa k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right].$
@@
$\leq\mathbb{E}\left[\left\|{\boldsymbol{\theta}^{0}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ch^{2}+C(1+\kappa)k^{\frac{1}{2}-\delta}+C(1+\kappa)k\sum_{\ell=1}^{n}\mathbb{E}\left[\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right].$ (3.41)
@@
$=:\frac{1}{2}\mathbb{E}\left[\left\|{\nabla\boldsymbol{\theta}^{0}}\right\|_{\mathbb{L}^{2}}^{2}\right]+I_{1}+I_{2}+\cdots+I_{13}.$ (3.44)
We will estimate each term on the last line. In the following, whenever appropriate, we always use the results of Proposition 2.2 and 3.5 without further mention. Let $\epsilon>0$. For the first term, by (2.16) and Young’s inequality we have
@@
$\leq Ce^{C\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right)+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right].$
@@
$\leq Cnh^{4}k^{-1}+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$
$\leq Ck+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right],$
where in the last step we used the assumption $h=O(k)$. Similarly, noting Proposition 2.2 we have
@@
$\leq Ce^{C\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right)+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right],$
@@
For the next term, we write $\boldsymbol{u}(t_{\ell})=\boldsymbol{u}(t_{\ell})-\boldsymbol{u}(t_{\ell-1})+\boldsymbol{u}(t_{\ell-1})$ and utilise the Hölder continuity in time of $\boldsymbol{u}$ to obtain
$\left\|{\boldsymbol{u}(t_{\ell})}\right\|_{\mathbb{L}^{\infty}}\leq k^{\frac{1}{2}-\delta}\left\|{\boldsymbol{u}}\right\|_{\mathcal{C}^{1/2-\delta}_{T}(\mathbb{H}^{2})}+\left\|{\boldsymbol{u}(t_{\ell-1})}\right\|_{\mathbb{L}^{\infty}}.$
To estimate $I_{6}$, we apply Young’s inequality, (2.10), (2), and (2.13). By Proposition 2.2 (Hölder continuity of $\boldsymbol{u}$) and the Sobolev embeddings, we obtain for any $\delta>0$,
@@
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ck^{1-\delta}$
@@
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+C(1+\kappa^{4})e^{C\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right),$ (3.46)
@@
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{4}}^{2}\right]+Ch^{4}\mathbb{E}\left[\left\|{\boldsymbol{H}}\right\|_{L^{\infty}_{T}(\mathbb{L}^{2})}^{4}+\left\|{\boldsymbol{u}}\right\|_{L^{\infty}_{T}(\mathbb{H}^{2})}^{4}\right]$
$\quad+Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left(k^{1-\delta}\left\|{\boldsymbol{H}}\right\|_{\mathcal{C}^{1/2-\delta}_{T}(\mathbb{L}^{2})}^{2}+\left\|{\boldsymbol{H}(t_{\ell-1})}\right\|_{\mathbb{L}^{2}}^{2}\right)\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{4}}^{2}\right]$
@@
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\Delta_{h}\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+C(1+\kappa)e^{C\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right),$
@@
where in the last step we also used (2.16) and (2.22), Proposition 3.5, and the Hölder continuity of $\boldsymbol{u}$. Next, we estimate $I_{8}$. To this end, noting (3.30), (2.5b), and (3.1), we write
$\boldsymbol{\eta}^{\ell}+\boldsymbol{\xi}^{\ell}$ $=\boldsymbol{H}(t_{\ell})-\boldsymbol{H}_{h}^{n}$
$=\Delta\boldsymbol{u}(t_{\ell})-\Delta_{h}\boldsymbol{u}_{h}^{\ell}+f_{R}(\boldsymbol{u}(t_{\ell}))-\Pi_{h}f_{R}(\boldsymbol{u}_{h}^{\ell-1})$
$=(I-\Pi_{h})\Delta\boldsymbol{u}(t_{\ell})+\Delta_{h}\boldsymbol{\theta}^{\ell}+(I-\Pi_{h})f_{R}(\boldsymbol{u}(t_{\ell}))+\Pi_{h}\big(f_{R}(\boldsymbol{u}(t_{\ell}))-f_{R}(\boldsymbol{u}_{h}^{\ell-1})\big),$
where in the last step we used the fact that $\Delta_{h}\mathcal{R}_{h}=\Pi_{h}\Delta$. It follows that
@@
$\left|{I_{8}}\right|$ $\leq\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left|{\left\langle\boldsymbol{u}_{h}^{\ell}\times\left((I-\Pi_{h})\Delta\boldsymbol{u}(t_{\ell})\right),\Delta_{h}\boldsymbol{\theta}^{\ell}\right\rangle}\right|\mathrm{d}s\right]$
@@
$\quad+\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left|{\left\langle\boldsymbol{u}_{h}^{\ell}\times\left((I-\Pi_{h})f_{R}(\boldsymbol{u}(t_{\ell}))\right),\Delta_{h}\boldsymbol{\theta}^{\ell}\right\rangle}\right|\mathrm{d}s\right]$